OpenStudy (anonymous):

if an ellipitical whisperingn room has a height of 30 ft and a width of 100ft, where should two people stand if they would like to whisper back and forth and be hard.

OpenStudy (turingtest):

So that basically is just asking what the foci of your ellipse are

OpenStudy (anonymous):

At the focii

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

I forget the formula for the foci, let me look it up

OpenStudy (anonymous):

its A2-B2=C2

OpenStudy (turingtest):

great, so what is the issue here?

OpenStudy (anonymous):

umm welll so would i do 100-30

OpenStudy (turingtest):

you understand that the formula is\[b^2-a^2=c^2\]so you have to square a and b...\[c=\sqrt{b^2-a^2}\]and the foci are\[(-c,0),(c,0)\]

OpenStudy (anonymous):

nooo its a^2-b^2=C^2

OpenStudy (anonymous):

depends on your major axis whether you add c to the x-coordinate of the center or y-coordinate

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

ok so was i right when i said 100-30=70

OpenStudy (anonymous):

a^2 is not 100, b^2 is not 30

OpenStudy (anonymous):

they gave you the width and height of the ellipse, not a and b

OpenStudy (anonymous):

okkk so how do i find a and b thats my issue

OpenStudy (anonymous):

They are half the width and height

OpenStudy (anonymous):

50-15

OpenStudy (anonymous):

i know that part

OpenStudy (anonymous):

so for horizontal major axis, a>b a = 1/2 width b= 1/2 height

OpenStudy (anonymous):

a=50 b=15

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now c^2=a^2-b^2

OpenStudy (anonymous):

solve for c, the focii are at (c,0) and (-c,0)

OpenStudy (anonymous):

ok so 50^2-15^2

OpenStudy (anonymous):

\[c=\sqrt{50^2-15^2}\]

OpenStudy (anonymous):

50^2=2500 15^2=225 2275 square root of that is 47.69

OpenStudy (anonymous):

\[5\sqrt{91} \approx 47.70\]

OpenStudy (anonymous):

ok same thing so since we know the foci so should people stand 47ft. away in order for them to whisper back and forth

OpenStudy (anonymous):

no, the focii are 47.7 ft from the center (0,0), so they'd stand twice as far away as 47.7, but be standing at the focii

OpenStudy (anonymous):

person a would be standing at (5sqrt(91),0), person b would be standing at (-5sqrt(91),0)

OpenStudy (anonymous):

ohh ok so do you think i would have to graph this

OpenStudy (anonymous):

|5sqrt(91)+5sqrt(91)|=10sqrt(91)

OpenStudy (anonymous):

does it tell you to graph it?

OpenStudy (anonymous):

i wouldn't

OpenStudy (anonymous):

ohh okay tynkss soo ,muchh

OpenStudy (anonymous):

np